# Evidence of Memory Effects in the Dynamics of Two-Level System Defect Ensembles Using Broadband, Cryogenic Transient Dielectric Spectroscopy

### v1

@@ III Results
As described in Sec. II, our waveguide enables a modular characterization of the driven and transient dielectric response of TLS defect ensembles in various samples. The samples considered in this study started with a 50.8 mm diameter, 432 $\mu$m thick, wafer of ultra-high-purity HEMEX sapphire from Crystal Systems, which has been measured to have bulk loss tangents of the order $10^{-7}$ [13]. We spin coat the samples with photoresist and dice them into rectangular chips with 27.5 $\times$ 5.5 mm dimensions. We clean the samples with solvents following a standard procedure described in Appendix B and mount them in the waveguide (Fig. 2a), which is cooled to 10 mK in a dilution refrigerator.
@@
Figure 3 summarizes the dielectric spectroscopy results obtained for different materials, alongside a room temperature control experiment of bare sapphire samples that serves as a benchmark. Given that the TLS defects observed in our experiments have frequencies between 3 and 5 GHz, small temperature changes can significantly impact their thermal occupation, a key signature that has been explored in previous work [48, 49]. Furthermore, increases in temperature cause thermal occupation of phonon modes, which can cause incoherent shaking of the double-well potential of the TLS defects, likely broadening or destroying any spectral features. As motivated in Sec. II, linear response theory suggests that the spectral dips of the driven dielectric response should be corroborated by the longer-lived coherence times in the transient region, which is saturated at room temperature (Fig. 3a, ii).
In contrast, when the same sample in Fig. 3a is cooled down to 10 mK, we see an apparent narrowing of the spectral dips in the driven dielectric spectroscopy (Fig. 3b,i) and a corresponding increase in the coherence of the transient dielectric response (Fig. 3b,ii). Appendix C provides a more detailed demonstration of this effect (using the sample in Fig. 3c) by zooming into a narrower frequency range and varying the pulse lengths. Most emission locations align with the absorption minima during the drive. As the pulse length increases, the absorption features deepen, and the associated emission peaks exhibit increased sharpness, providing clear signatures of coherent energy storage and re-emission by the TLS defect ensemble. In Fig. 3 b, we conjecture that these defects arise from residual surface contamination due to leftover photoresist, adventitious carbon, surface damage, dicing damage, or the waveguide and antenna surfaces. In future work, we intend to explore further surface treatments such as etching, annealing, and passivation to reduce the measured transient dielectric response of bare sapphire and silicon wafers [50, 51, 52].
@@
Today, most Josephson junctions are fabricated using various versions of double-angle evaporation, where the insulating Josephson junction layer is formed from a native oxide grown on deposited aluminum. However, all regions of the chips that contain aluminum will inadvertently grow such native oxides, and therefore, most superconducting circuits inevitably contain large areas of aluminum oxide. To emulate the native oxide in Josephson junctions and aluminum metal layers, we grow a thin, 2 nm aluminum oxide layer in an atomic layer deposition tool on a bare sapphire wafer, like those measured in Fig. 3a-b (see Appendix B). The results of these dielectric spectroscopy measurements are shown in Fig. 3c. Here, we see obvious signatures of TLS defect ensembles, primarily in the transient region. The transient response exhibits pronounced collapse and revivals due to the interplay between external driving and TLS-defect ensemble interactions, clearly indicating memory effects in the environment [20, 53]. Furthermore, we want to highlight that the deposited 2 nm of aluminum oxide is approximately $5\times 10^{-6}$ times the volume of the substrate, but is composed of the same atomic species. These findings demonstrate the importance of material morphology in the existence and behavior of TLS defects and motivates future work on passivating metallic layers [54] and developing crystalline insulating layers for Josephson junctions. This work also suggests why tantalum, which hosts a thin, stoichiometric native oxide, is an advantageous material choice for high-coherence superconducting circuits [50, 26].
@@
Finally, we also characterize the dielectric response of 1-4 $\mu$m of photoresist on sapphire, as shown in Fig. 3d. This investigation is especially relevant because most resonators and waveguides leverage photolithography, meaning residual photoresist may be left on substrates if improperly removed. These photoresists are complex polymer chains that likely contain a high density of TLS defects [55, 4] (a detailed description of sample preparation is included in Appendix B). By understanding the dielectric properties of the resist layer, we can better assess its impact on junction performance and guide fabrication and device design optimizations [42]. Therefore, in Fig. 3d, we observe more transient response, consistent with our hypothesis that photoresist has many TLS defects.
@@
The technique presented in this manuscript also enables future longitudinal studies where the sample can be characterized at each step of the fabrication process, and it envisions active mitigation strategies for these defects in the next generation of superconducting qubit devices. It is essential to highlight that in the data presented in Fig. 3, we observe that the transient emission processes have lifetimes on the order of $100$ ns. Previous work on so-called dielectric echoes in JTWPA experiments [39, 40] measured lifetimes of  $\approx 1$ $\mu$s. Here, since the waveguide provides a large density of states and probes large ensembles, the lifetimes of these TLS defects may be reduced due to Purcell enhancement and stimulated emission. While the lifetimes extracted in these measurements are relatively short compared to state-of-the-art coherence times for superconducting qubits, the observed transient response can still play a prominent role in larger quantum circuits where gates are sequentially implemented. This effect may partially explain why large quantum circuits have overall fidelities lower than a naive calculation based on single and two-qubit gate fidelities extracted using randomized benchmarking [56].
Cryogenic Dielectric Spectroscopy of the Same Sample, For Different Cooldowns
@@
Figure 5: Transient dielectric response of Shipley 1813 photoresist on sapphire, same data as Fig. 3d. a-ii, Logarithmic magnitude of the transient dielectric response. b-ii, Magnitude of Fast Fourier transform (FFT) of the logarithmic amplitude. c, Log of the two-time correlation function g(2) map of the transient spectrum, defined by equation 15. d-ii, $\chi^{\prime\prime}$ computed from g(2) using equation 14. We highlight transitions from positive to negative $\chi^{\prime\prime}$ with the cyan contour in d-ii. a-d,i show linecuts at $\omega/2\pi=3.322$ GHz (dashed line in ii), highlighting distinct memory effects in the form of collapse and revivals in the ringdown (a), sharp spectral features in the FFT (b), non-exponential decay, bunching, and oscillations indicative of non-Markovian dynamics (c), non-monotonic and negative $\chi^{\prime\prime}$ (d). The negative regions are consistent with the bright FFT features in (b). It should be noted that $\chi^{\prime\prime}$ oscillates around zero near $\omega/2\pi=3$ GHz primarily due to weak signals and not necessarily indicating memory effects. A short horizontal white line is added to indicate pulse bandwidth $\Delta$f.
Much like point defects in crystals, which spatially migrate at high temperatures, it is well-understood that thermal cycling of TLS defects, even up to a few K, can cause them to move considerably in frequency and potentially in physical location. To corroborate this signature of TLS defects, we perform dielectric spectroscopy experiments 15 days apart on the same sample subject to a thermal cycle to 300 K in Fig. 4. As expected, we see a significantly different driven and transient dielectric response in Fig. 4a-b, highlighted by the difference plot in Fig. 4c. In future work, we intend to perform systematic studies of the temperature dependence closer to 10 mK to observe the thermal occupation and ultimate dissolutions of TLS defects. Such studies may provide insights into the observed ‘universal’ dielectric response, and potentially probe the transition of quantum TLS defects to the broad classical dielectric response [57, 58].
@@
In this section, we provide a more in-depth analysis of the transient dielectric response from the sample in Fig. 3d containing photoresist on sapphire, which is shown in Fig. 5. We start by plotting the transient dielectric response in Fig. 5a,ii showing long-timescale oscillatory collapse and revival behavior of the emitted signal, which indicates information backflow and memory effects [59, 60]. This can be seen clearly in Fig. 5a,i, which displays the response at a single frequency of $\omega/2\pi=3.322$ GHz. To quantify the frequency components of the oscillations observed in the time traces in Fig. 5a, we perform Fast Fourier Transforms (FFT) of all the data traces to generate Fig. 5.b,ii. Again, a particular line cut at the same frequency is shown in Fig. 5b,i. The FFT clearly shows the emergence of distinct V-shaped frequency components and horizontal bands across parts of the FFT spectrum (e.g. at frequency range 3.65-4.0 GHz, and 4.25-4.65 GHz). These V-shapes can be attributed to the off-resonant emission of the dominant eigenfrequencies of the TLS defect ensembles, which are prominent due to the broadband nature of the waveguide [61].
@@
where $\mathcal{I}(t)\propto A^{2}(t)$ is the intensity of the transmitted pulse, and plot it in Fig. 5c. $g^{(2)}(\tau^{\prime})$ provides us with a map of temporal correlations and displays evident non-exponential, oscillatory decay, which is a hallmark signature of memory effects and non-Makovian dynamics of the TLS defect ensemble. Finally, we utilize the computed $g^{(2)}(\tau^{\prime})$ to extract $\chi^{\prime\prime}(\omega)$ using Eq. 14, which we show in Fig. 5d. Here, consistent with the $g^{(2)}(\tau^{\prime})$, we observed the non-monotonic and negative $\chi^{\prime\prime}(\omega)$, accentuated by the cyan contours in Fig. 5d, further supporting the claim of memory effects and re-emission in the TLS defect ensemble. The cyan contour highlights the negative areas of $\chi^{\prime\prime}(\omega)$, which indicate physics beyond linear response, exhibiting partial overlap with the aforementioned V-shapes in some regions of the FFT [63, 64].
@@ IV Conclusions and Outlook
This work introduces cryogenic transient dielectric spectroscopy, a novel technique to probe TLS defects that supports fast, modular mounting of various samples. Our measured cryogenic dielectric spectroscopy experiments reveal distinct transient features and memory effects as demonstrated by the $g^{(2)}$ and $\chi^{\prime\prime}$ plots in Fig. 5c,d, respectively. Such memory effects may indicate that the dynamics are non-Markovian, which could create problems for quantum error correction if manifested in quantum processors [65, 66, 53]. We believe that much of the memory effects arise from the interplay between coherent TLS defects and driving in a broadband environment. By analyzing the spatial and temporal correlations of large TLS defect ensembles in isolated materials, we take initial steps toward uncovering the atomistic structure of TLS defects. Next, we aim to introduce an orthogonal SMA port to rotate the polarization of the drive and perform the same dielectric spectroscopy experiments, subject to different drive strengths. By integrating over different drive polarizations, we can better understand the frequency distribution and dipole moments of the TLS defects in various samples.
While this work has focused on materials for superconducting qubits, we highlight that cryogenic transient dielectric spectroscopy is a more general technique that provides information about the atomistic structure of various dielectric materials. The response is also not exclusively driven by TLS defects, and we may be able to measure point defects as well. For instance, transient dielectric spectroscopy of silicon (diamond) samples may allow us to measure silicon (nitrogen) vacancies without needing optical interrogation. In conclusion, we have demonstrated a novel spectroscopy tool based on a broadband 3D waveguide that enables global interrogation of TLS defect ensembles. The approach is entirely modular, and we demonstrate that this can be leveraged for longitudinal devices that can help develop causal relationships between TLS defect ensemble properties and material processing steps.
@@ A.1 Waveguide Design
Figure 7 provides detailed dimensions of our SMA to WR-229 3D rectangular waveguide adapter. In this work, we leverage Eq. 19 to achieve the desired 2.8 to 6 GHz frequency band and adapt the perfect rectangular waveguide designs to make the manufacturing easier. In particular, our design features an inner radius of the 3D waveguide designed to use a 3/8 inch end mill. M2.5 screws are used to mount the 4.1 mm SMA pins that integrate into the designs shown in Fig. 7.
@@ Appendix B Sample Processing
We start with high-grade crystalline 0.5 mm-thick sapphire wafers from Crystal Systems, which are then diced into 27.5 x 5.5 mm rectangular strips, the standard substrate size for hosting 3D transmons. The samples are initially covered in resist AZ 1518 to protect them from debris during dicing (likely another source of TLS defects). We soak the chips in acetone at 70 C overnight and sonicate them in acetone before transferring them to IPA and drying with compressed air.
@@
Figure 8: Dielectric response of 2nm aluminum oxide under different pulse lengths. a, 20 ns pulse. b, 50 ns pulse. c, 200 ns pulse. As the pulse length increases, the bandwidth (marked by the short vertical white lines) of the pulse decreases, resulting in sharper emissions. Absorption in the pulsing region often corresponds to an emission around the same frequency, after the pulse is turned off.
Appendix C Dielectric Response under Different Drive Lengths
We examine the driven and transient dielectric response under drive pulses of varying lengths in Fig. 8, featuring 20 ns (a), 50 ns (b), and 200 ns (c) pulses. To clearly display subtle features, we restrict the analysis to a narrower frequency window of 3.9-4.4 GHz.
Increasing the drive length improves spectral resolution and strengthens on-resonance driving on TLS defects. This manifests as sharper and deeper absorption dips during the pulse, due to increased spectral selectivity as the drive couples more effectively to resonant TLS defects. After the drive is turned off, a similar effect is seen in the transient emission, where we observe increased sharpness and prolonged lifetimes in correspondence to the reduced bandwidth of the drive. This result is consistent with both our conceptual model and numerical simulations (to be presented elsewhere). Moreover, even without controlled access to physical parameters of the TLS defects, it is encouraging to observe changes in TLS dynamics as a result of different pulses. These findings support the feasibility of pulse-based coherent control of TLS defect ensembles.

### v2

@@ III Results
As described in Sec. II, our waveguide enables a modular characterization of the driven and transient dielectric response of ensembles of TLS defects in various samples. The samples considered in this study started with a 50.8 mm diameter, 432 $\mu$m thick, wafer of ultra-high-purity HEMEX sapphire from Crystal Systems, which has been measured to have bulk loss tangents of the order $10^{-8}$ using the dielectric dipper technique [13]. We spin coat the samples with photoresist and dice them into rectangular chips with 27.5 $\times$ 5.5 mm dimensions. We clean the samples with solvents following a standard procedure described in Appendix  B and mount them in the waveguide (Fig. 2a).
@@
Figure 3 summarizes the dielectric spectroscopy results obtained for different materials, alongside two room temperature control experiments of empty waveguide (Fig. 3a) and bare sapphire samples (Fig. 3b) that serves as a benchmark. Given that the TLS defects observed in our experiments have frequencies between 3 and 5 GHz, small temperature changes can significantly impact their thermal occupation, a key signature that has been explored in previous work [49, 50]. Furthermore, increases in temperature cause thermal occupation of phonon modes, which can lead to incoherent shaking of the double-well potential governing the TLS defects, likely broadening or destroying their spectral features. Our room temperature measurements (Fig. 3aii and bii) confirm this thermal saturation effect.
As motivated in Sec. II, when we cool down to cryogenic temperatures and remove this thermal effect, linear response theory suggests that the spectral dips of the driven dielectric response should be corroborated by the longer-lived coherence times in the transient region. This is consistent with our findings when we cool down the same setups in Fig. 3a and b to 10 mK, we see an apparent narrowing of the spectral dips in the driven dielectric spectroscopy and a corresponding increase in the coherence of the transient dielectric response (Fig. 3c and d). These features are clear signatures of coherent energy storage and re-emission by the ensembles of TLS defects, which we discuss further in the context of master equation and Floquet theory in the next section (with further investigations in Appendix  G). In Fig. 3c, we observe defect features in the absence of mounted samples, which we attribute to oxide layers present on the waveguide and antenna surfaces, and uncontrolled modes in the dilution fridge enclosure that couple to the waveguide. The signal in Fig. 3d likely originates from similar sources but may also include contributions from the solvent-cleaned sapphire samples in the waveguide, where residual surface contamination can result from leftover photoresist, adventitious carbon, or surface and dicing damages. In future work, we intend to remove these baseline contributions by retracting the samples in situ with a piezoelectric positioner, inspired by the dielectric dipper approach [13]. This calibration will enable further exploration of surface treatments such as etching, annealing, and passivation to reduce the measured transient dielectric response of bare sapphire and silicon wafers, which form the basis of all superconducting circuits [51, 52, 53].
@@
Today, most Josephson junctions are fabricated using various versions of double-angle evaporation, where the insulating Josephson junction layer is formed from a native oxide grown on deposited aluminum. However, all regions of the chips that contain aluminum will inadvertently grow such native oxides, and therefore, most superconducting circuits inevitably contain large areas of aluminum oxide. To emulate the native oxide in Josephson junctions and on aluminum metal layers, we grow a thin, 2 nm aluminum oxide layer in an atomic layer deposition tool on a bare sapphire wafer (Fig. 3b and d). The results of these dielectric spectroscopy measurements are shown in Fig. 3e. Here, we see clear signatures of ensembles of TLS defects, primarily in the transient region. The transient response exhibits pronounced collapse and revivals due to the interplay between external driving and TLS-defect ensemble interactions, clearly indicating memory effects [4, 54]. Furthermore, we want to highlight that the deposited 2 nm of aluminum oxide is approximately $5\times 10^{-6}$ times the volume of the substrate, but is composed of the same atomic species. These findings demonstrate the importance of material morphology in the existence and behavior of TLS defects and motivate future work on passivating metallic layers [55] and developing crystalline insulating layers for Josephson junctions. This work also suggests why tantalum, which hosts a thin, stoichiometric native oxide, is an advantageous material choice for high-coherence superconducting circuits [51, 25].
@@
Finally, we also characterize the dielectric response of 1-4 $\mu$m of photoresist on sapphire, as shown in Fig. 3f. This investigation is especially relevant because most resonators and waveguides leverage photolithography, meaning residual photoresist may be left on substrates if improperly removed. These photoresists are complex polymer chains that likely contain a high density of TLS defects [56, 57] (a detailed description of sample preparation is included in Appendix  B). By understanding the dielectric properties of the resist layer, we can better assess its impact on junction performance and guide fabrication and device design optimizations [43]. Therefore, in Fig. 3f, we observe more transient response, consistent with our hypothesis that photoresist has many TLS defects.
@@
The technique presented in this manuscript also enables future longitudinal studies where the sample can be characterized at each step of the fabrication process, and it envisions active mitigation strategies for these defects in the next generation of superconducting qubit devices. It is essential to highlight that in the data presented in Fig. 3, we observe that the transient emission processes have lifetimes on the order of $100$ ns. Previous work on so-called dielectric echoes in JTWPA experiments [40, 41] measured lifetimes of  $\approx 1$ $\mu$s. Here, since the waveguide provides a large density of states and probes large ensembles, the lifetimes of these TLS defects may be reduced due to Purcell enhancement and stimulated emission. While the lifetimes extracted in these measurements are relatively short compared to state-of-the-art coherence times for superconducting qubits, the observed transient response can still play a prominent role in larger quantum circuits where gates are sequentially implemented. This effect may partially explain why large quantum circuits have overall fidelities lower than a naïve calculation based on single and two-qubit gate fidelities extracted using randomized benchmarking [58].
Figure 4: Dielectric response of 2 nm aluminum oxide sample under different pulse lengths. a, 20 ns pulse. b, 50 ns pulse. c, 200 ns pulse. Similar to Fig. 3, subpanels ii show the driven and transient dielectric response, while subpanels i show the average of the driven dielectric response and the extracted lifetimes ($\tau$) of the transient dielectric response. As the pulse length increases, the bandwidth (marked by the short vertical white lines) of the pulse decreases, resulting in sharper emissions. Absorption in the pulsing region often corresponds to an emission around the same frequency, after the pulse is turned off.
Dielectric Response under Different Drive Lengths
In order to better examine the relationship between absorption in the driven dielectric response and the transient emission, we zoom into a narrower frequency range of 3.9-4.4 GHz and vary the pulse lengths in Fig. 4, featuring 20 ns (Fig. 4a), 50 ns (Fig. 4b), and 200 ns (Fig. 4c) pulses. Increasing the drive length improves the spectral resolution and reveals finer resonant features, which manifest themselves as sharper absorption dips during the pulse. After the drive is turned off, a similar effect is seen in the transient emission, where we observe sharper and more prolonged ringdowns. Without controlled access to the physical parameters of the TLS defects, it is encouraging to observe changes in the TLS dynamics as a result of different pulses. These findings support the feasibility of pulse-based coherent control of ensembles of TLS defects.
To explore the connection between spectral dips and the surrounding emission features, we performed numerical simulations of two interacting driven spins. The full pulse evolution was treated using a master equation, while the periodic portion of the drive was analyzed separately using Floquet formalism, as detailed in Appendix  G and illustrated in Fig. 12. During the pulse, the system is dressed by the external periodic drive, resulting in the emergence of sidebands. The corresponding Floquet quasi-energies show clear dips (Fig. 12b), and prolonged ring-downs within the vicinity of these dips, as seen in Fig. 12a. In Fig. 13, we qualitatively reproduce the sharpening of ring-downs for different pulse durations as observed in the experiment.
@@
Figure 5: Transient dielectric response of Shipley 1813 photoresist on sapphire, same data as Fig. 3d. aii, Logarithmic magnitude of the transient dielectric response. bii, Magnitude of Fast Fourier transform (FFT) of the logarithmic amplitude. cii, Log of the two-time correlation function g(2) map of the transient spectrum, defined by Eq. 15. dii, $\chi^{\prime\prime}$ computed from g(2) using Eq. 14. We highlight transitions from positive to negative $\chi^{\prime\prime}$ with the cyan contour in dii. Subpanels i show linecuts at $\omega/2\pi=3.322$ GHz (dashed line in ii), highlighting distinct memory effects in the form of collapse and revivals in the ringdown (a), sharp spectral features in the FFT (b), non-exponential decay, bunching, and oscillations indicative of non-Markovian dynamics (c), and non-monotonic and negative $\chi^{\prime\prime}$ (d). The negative regions are consistent with the bright FFT features in bii. It should be noted that $\chi^{\prime\prime}$ oscillates around zero near $\omega/2\pi=3$ GHz primarily due to weak signals and not necessarily indicating memory effects. A short horizontal white line is added to indicate pulse bandwidth $\Delta$f.
@@
In this section, we provide a more in-depth analysis of the transient dielectric response from the sample in Fig. 3d containing photoresist on sapphire, which is shown in Fig. 5. We start by plotting the transient dielectric response in Fig. 5aii showing long-timescale oscillatory collapse and revival behavior of the emitted signal, which indicates information backflow and memory effects [59, 60]. This can be seen clearly in Fig. 5ai, which displays the response at a single frequency of $\omega/2\pi=3.322$ GHz. To quantify the frequency components of the oscillations observed in the time traces in Fig. 5a, we perform a FFT of all the data traces to generate Fig. 5bii. Again, a particular line cut at the same frequency is shown in Fig. 5bi. The FFT shows clear V-shaped features and horizontal bands, particularly in the 3.65–4.0 GHz and 4.25–4.65 GHz ranges. These V-shapes arise from off-resonant emission by dominant eigenfrequencies of the ensembles of TLS defects, made visible by the broadband response of the waveguide [61]. Numerical simulations of a driven, interacting four-spin system (Appendix  G, Fig. 13) confirm this behavior, namely the appearance of V-shaped arcs in the FFT, which sharpen with longer pulse durations, and converge toward the bare eigenmode frequencies.
@@
where $\mathcal{I}(t)\propto A^{2}(t)$ is the intensity of the transmitted pulse, and plot it in Fig. 5c. $g^{(2)}(\tau^{\prime})$ provides us with a map of temporal correlations and displays evident non-exponential, oscillatory decay, which is a hallmark signature of memory effects or non-Makovian dynamics of the ensembles of TLS defects. Finally, we utilize the computed $g^{(2)}(\tau^{\prime})$ to extract $\chi^{\prime\prime}(\omega)$ using Eq. 14, which we show in Fig. 5d. Here, consistent with $g^{(2)}(\tau^{\prime})$, we observe non-monotonic and negative $\chi^{\prime\prime}(\omega)$, highlighted by the cyan contours at zero in Fig. 5d, exhibiting partial overlap with the aforementioned V-shapes in some regions of the FFT. This behavior indicates physics beyond linear response and further supports the presence of memory effects and re-emission in the ensembles of TLS defects [63, 64].
@@ IV Conclusions and Outlook
This work introduces BCTDS, a technique to probe TLS defects that supports fast and modular mounting of various samples, without invoking narrow-band probes like resonators or qubits. Our measured cryogenic dielectric spectroscopy experiments reveal distinct transient features and memory effects as demonstrated by the $g^{(2)}$ and $\chi^{\prime\prime}$ plots in Fig. 5c,d, respectively. Such memory effects may indicate that the dynamics are non-Markovian, which may be detrimental for quantum error correction if manifested in quantum processors [65, 66, 54]. We believe that much of the memory effects arise from the interplay between coherent TLS defect interactions and driving and measuring in a broadband environment. By analyzing the frequency and temporal correlations of large ensembles of TLS defects in isolated materials, we take initial steps toward uncovering the atomistic structure of TLS defects. Next, we aim to introduce an orthogonal SMA port to rotate the polarization of the drive and perform the same dielectric spectroscopy experiments, subject to different drive strengths. By integrating over different drive polarizations, we can better understand the frequency distribution and dipole moments of the TLS defects in various samples.
While this work has focused on materials for superconducting qubits, we highlight that BCTDS is a more general technique that provides information about the loss of various dielectric materials. This study is also not exclusively applicable to TLS defects; for instance, we can use BCTDS to measure vacancies and boron dopants in silicon and nitrogen vacancies in diamond directly, without the need for optical interrogation [67]. In conclusion, we have demonstrated a novel spectroscopy tool based on a broadband 3D waveguide that enables global interrogation of ensembles of TLS defects. The approach is entirely modular, and we demonstrate that this can be leveraged for longitudinal devices that can help develop causal relationships between properties of ensembles of TLS defects and material processing steps.
@@ A.1 Waveguide Design
Figure 7 provides detailed dimensions of our SMA to WR-229 3D rectangular waveguide adapter. In this work, we leverage Eq. 19 to achieve the desired 3 to 6 GHz frequency band and adapt the perfect rectangular waveguide designs to make the manufacturing easier. In particular, our design features an inner radius of the 3D waveguide designed to use a 3/8 inch end mill. M2.5 screws are used to mount the 4.1 mm SMA pins that integrate into the designs shown in Fig. 7.
@@ Appendix B Sample Processing
We start with high-grade crystalline 0.5 mm-thick sapphire wafers from Crystal Systems, which are then diced into 27.5 x 5.5 mm rectangular strips, the standard substrate size for hosting 3D transmons. The samples are initially covered in resist AZ 1518 to protect them from debris during dicing (likely another source of TLS defects). We soak the chips in acetone at 70 C overnight and sonicate them in acetone before transferring them to IPA and drying with compressed nitrogen gas.
@@
Appendix C Cryogenic Dielectric Spectroscopy of the Same Sample, For Different Cooldowns
Two-level system (TLS) defects are known to shift in frequency and spatial location after thermal cycling [68], due to rearrangements in the disordered environment that reshape the local energy landscape. To corroborate this signature, we perform dielectric spectroscopy experiments 15 days apart on the same sample subject to a thermal cycle to 300 K in Fig. 8. As expected, we see a significantly different driven and transient dielectric response in Fig. 8a,b, highlighted by the difference plot in Fig. 8c. In future work, we intend to perform systematic studies of the temperature dependence closer to 10 mK. Such studies may provide insight into the observed ‘universal dielectric response’, and potentially probe the transition of quantum TLS defects to the broad classical dielectric response [69, 70].
Figure 9: Dielectric response of Shipley 1813 photoresist on sapphire measured in transmission and reflection. We used the same samples as in the thermal-cycling experiment in Fig. 8, but we add a cryogenic circulator (QCY-G0400801) to the waveguide input to measure both the transmitted (a) and reflected (b) signals. We also include a difference plot (c), which shows a similarity in response from the transient region.
Appendix D Dielectric Response in Transmission and Reflection
For ideal, lossless systems, transmission and reflection contain equivalent information as incident power must be conserved. In our experiment, where dissipation is present, reflection provides complementary information about dielectric loss during the pulsing region. We make measurements on the same samples as in Appendix  C but place a cryogenic circulator at the waveguide input to measure both the transmission (Fig. 9a) and reflection (Fig. 9b) signals. In the transient region, emission from the dielectric samples is detected almost equally in the transmission and reflection channels, as the samples are positioned equidistant from the two antenna pins, and the readout circuits are identical. As a result, the collected signal shows minimal difference, as shown in Fig. 9c. A short vertical white line is added to indicate pulse bandwidth $\Delta$f.